# Non-unitary enhanced transfer efficiency in quantum walk search on complex networks

### v2

@@ B.2 Ring-lattice transition
As the results on the cycle and the complete graphs are completely different, we use the Ring-lattice graph model to progressively transform a cycle into a complete graph. A ring-lattice graph also known as a $k$-cycle is the basis of Watts and Strogatz model widely known as small-world networks [67]. It consists of a cycle graph where each vertex is connected to its $k$ nearest neighbors. Therefore when $k=2$ we recover a cycle graph and the maximum value of $k$ generates a complete graph. As an illustration we show the transition from the cycle to the complete graph of size $N=8$ in Fig. 9. We run the SQWS on the Ring-lattice graph of size $N=32$ for 16 differents values of $k$, the cycle graph is obtained for $k=2$ and the complete graph for $k=32$. We present the results in Fig. 11. As $k$ increases, the eccentricity of the marked vertex decreases as the graph’s connections increase, leading to an increase in its centrality as shown in Fig. 10. We observe that the hybrid regime outperforms the purely quantum one only for the cycle graph, i.e. $k=2$. A slight increase in connectivity allows the quantum regime to gain the upper hand over the hybrid, as we can see that from $k=4$ the hybrid regime is no more efficient than the purely quantum one for low values of $\gamma$. We then note that as the value of $k$ increases, the low-noise hybrid regime, i.e. low value of $\omega$, is less efficient than a noisier hybrid. This feature was clearly visible for the complete graph $K_{64}$ in Fig. 2.
@@
Figure 11: Stochastic Quantum Walk Search (SQWS) performances with different values of $\gamma\in[0.05,50]$ on Ring-lattice graphs of size $N=32$. The cycle graph graph is recovered for $k=2$ and the complete graph for $k=32$. An increase in the value of $k$ corresponds to increasing the average connection of each vertex with its $k$-nearest neighbors. The time evolution is set to $t=320$ units of time. The quantum walk dynamics is recovered for $\omega=0$, the classical random walk for $\omega=1$, and a linear combination of the two when $\omega\in]0,1[$.

### v3

@@ B.2 Ring-lattice transition
As the results on the cycle and the complete graphs are completely different, we use the Ring-lattice graph model to progressively transform a cycle into a complete graph. A ring-lattice graph also known as a $k$-cycle is the basis of Watts and Strogatz model widely known as small-world networks [67]. It consists of a cycle graph where each vertex is connected to its $k$ nearest neighbors. Therefore when $k=2$ we recover a cycle graph and the maximum value of $k$ generates a complete graph. As an illustration we show the transition from the cycle to the complete graph of size $N=8$ in Fig. 11. We run the SQWS on the Ring-lattice graph of size $N=32$ for 16 differents values of $k$, the cycle graph is obtained for $k=2$ and the complete graph for $k=32$. We present the results in Fig. 13. As $k$ increases, the eccentricity of the marked vertex decreases as the graph’s connections increase, leading to an increase in its centrality as shown in Fig. 12. We observe that initially, with the graph cycle when $k=2$, the regime around $\omega=0.1$ is the most efficient of all, and the quantum regime does not perform well. Then, as the graph density increases, the other regimes perform better. Thus, from $k=6$ onwards, the $\omega=0.1$ regime is the least efficient for $\gamma<1$. With increased connectivity, we also see that the quantum regime supports higher values of $\gamma$ to maintain performance above 80%. We note that the results seem to stabilise and are no longer sensitive to increases in connectivity from $k=10$ onwards, which corresponds roughly to the point at which the graph’s eccentricity no longer really decreases with increases in average connectivity, as shown in Fig. 12.
@@
Figure 13: Stochastic Quantum Walk Search (SQWS) performances with different values of $\gamma\in[0,30]$ on Ring-lattice graphs of size $N=32$. The cycle graph is recovered for $k=2$ and the complete graph for $k=32$. An increase in the value of $k$ corresponds to increasing the average connection of each vertex with its $k$-nearest neighbors. The time evolution is set to $t=320$ units of time. The quantum walk dynamics is recovered for $\omega=0$, the classical random walk for $\omega=1$, and a linear combination of the two when $\omega\in]0,1[$.